THE QUESTIONS ARE: to calculate the magnitude of t

THE QUESTIONS ARE:

to calculate the magnitude of the acceleration

AND

what would the magnitude of the acceleration be if the friction is 20.0N

Thank you!
Answer #1

acceleration a = Fnet/m = (95-75-12)/28 = 0.2857 m/s^2

Magnitude of acceleration if friction is 20N,   a = 95-75-20)/28 = 0 m/s^2 Answer